# Expressivity Limits and Trainability Guarantees in Quantum Walk-based Optimization

### v1

@@ 3.1 Bounding the DLA Dimension for QWOA
We provide an upper bound on the DLA dimension for QWOA that scales quadratically in $m$. Recall that $m$ is the number of distinct cost values over the feasible solutions, or equivalently, the number of distinct eigenvalues of $H_{C}$ restricted to the feasible subspace. The key idea behind the proof is the following. Since QWOA is invariant under permutations of basis states, solutions with the same cost can be grouped together in the problem Hamiltonian $H_{C}$. Due to the diagonal structure of the problem Hamiltonian $H_{C}$ and the uniform structure of the mixing Hamiltonian $H_{M}$, the nested commutators of the QWOA generators will share the same partition structure of $m^{2}$ blocks, where each block has constant entries. This means that the degree of freedom in the DLA space generated by the QWOA ansatz is bounded by $\mathcal{O}(m^{2})$.
@@
For any instance of a combinatorial optimization problem, the DLA dimension of QWOA satisfies $\operatorname{dim}(\mathfrak{g}_{\operatorname{QWOA}})\leq m^{2}+1$.
@@
We first consider the unconstrained problems case. Let $x_{1},\dots,x_{m}$ be the possible $m$ distinct costs for a given instance of an unconstrained optimization problem. The problem Hamiltonian is diagonal in the computational basis and can be rewritten in the block diagonal form as
$H_{C}=\begin{bmatrix}x_{1}I&\bar{0}&\cdots&\bar{0}\\ \bar{0}&x_{2}I&\cdots&\bar{0}\\ \vdots&\vdots&\ddots&\vdots\\ \bar{0}&\bar{0}&\cdots&x_{m}I\\ \end{bmatrix}$, (12)
where $I$ is the identity matrix with the dimension equal to the multiplicity of its respective cost and $\bar{x}$ denotes a matrix with all elements equal to $x$. We call such a partitioning of $H_{C}$ into blocks a block partitioning pattern.
Note that, up to a global phase, the mixing operator $e^{-it_{l}H_{M}}$ is equivalent to the operator $e^{-it_{l}J}$, where $J=K_{|S|}+I$ (all ones matrix). Therefore, we can set $H_{M}=J$ without loss of generality and rewrite $H_{M}$ using the block partitioning pattern as
@@
Now, let $\Lambda$ be the set of matrices with the same block partitioning pattern as $H_{C}$ and $H_{M}$ such that all elements in any block are the same. In this case, note that $H_{M}\in\Lambda$, but (in general) $H_{C}\notin\Lambda$. From the definition of matrix multiplication, it follows that $H_{C},A]\in\Lambda$ and $A,B]\in\Lambda$ for $A,B\in\Lambda$. Consequently, the DLA of QWOA is generated by $H_{C}$ and the elements of $\Lambda$. It remains to determine the size of $\Lambda$. Notice that for any matrix in $\Lambda$, there are at most $m^{2}$ possible configurations. This yields $|\Lambda|=m^{2}$. Since the elements of $\Lambda$ and $H_{C}$ form the basis of the DLA, we have $\operatorname{dim}(\mathfrak{g}_{\operatorname{QWOA}})\leq m^{2}+1$, establishing the unconstrained case.
@@ 3.2 Overparametrization in QWOA for NPO-PB problem
We now present two remarks regarding item 1 of Theorem 2. While these remarks are stated in the context of item 1, analogous reasoning applies to item 2 by replacing $\mathsf{BPPO}$ and $\mathsf{PO}$ with $\mathsf{BP\text{-}APX}$ and $\mathsf{APX}$, respectively. Firstly, the inverse of item 1 of Theorem 2 does not hold, i.e, $\mathcal{P}\in\textsf{BPPO}$ does not necessarily imply that the optimal QWOA depth is polynomially bounded in the input size. Indeed, the fact that a problem belongs to BPPO does not guarantee that it can be efficiently solved by random sampling, which is the baseline performance that QWOA quadratically improves upon. Secondly, unless $\mathsf{PO}=\mathsf{BPPO}$, we cannot state that $\mathcal{P}\notin\mathsf{PO}$ implies that $p^{*}_{\mathcal{P}}\notin\mathcal{O}(\operatorname{poly}(s))$. Whether there exists a problem in $\textsf{BPP}-\textsf{PO}$ that QWOA can still solve with polynomial depth remains an open question.
@@
3.3 The Barren Plateau Problem
By combining Corollary 1, which polynomially bounds the DLA dimension for $\mathsf{NPO}\text{-}\mathsf{PB}$ problems, with Eq. (10), which relates the DLA dimension to the variance of the loss function for QWOA, we get the following theorem.
Theorem 4 (Absence of BPs for QWOA).
[…]
and BPs do not exist in the loss function of QWOA for $\mathcal{P}$.
Proof.
First, note that
[…]
since projectors are trace-preserving. Then, we can write Eq. (10)) as follows.
[…]
From Corollary 1, we have
[…]
establishing Eq. (14). Since the variance of the loss function does not vanish exponentially with $s$, there are no BPs in the loss function of QWOA for $\mathcal{P}$. ∎
Theorem 4 reinforces and generalizes the argument of Headley and Wilhelm [28], who showed that QWOA avoids barren plateaus if the probability distribution associated with the solution space converges asymptotically to a fixed distribution. Their argument relies on the fact that, under such convergence, the distribution — and consequently the loss function — of QWOA becomes independent of the input size, thereby preventing vanishing gradients in the optimization landscape. On the other hand, our result is more general since it holds for any $\mathsf{NPO}\text{-}\mathsf{PB}$ problem, without any additional assumptions on the cost function.
@@ 4.1 Unstructured Search
$H_{1}=H_{C}=\begin{pmatrix}I&\overline{0}\\ \overline{0}&\overline{0}\\ \end{pmatrix},\quad H_{2}=H_{M}=\begin{pmatrix}\overline{1}&\overline{1}\\ \overline{1}&\overline{1}\\ \end{pmatrix}$. (19)
@@
$H_{3}=\begin{pmatrix}\overline{0}&\overline{1}\\ \overline{-1}&\overline{0}\\ \end{pmatrix},\quad H_{4}=\begin{pmatrix}\overline{0}&\overline{1}\\ \overline{1}&\overline{0}\end{pmatrix},\quad H_{5}=\begin{pmatrix}\overline{2(|M|-|S)}&\overline{2|M|-|S|}\\ \overline{2|M|-|S|}&\overline{2|M|}\\ \end{pmatrix}$. (20)
If $|M|>1$, one can show that $\{H_{1},H_{2},H_{3},H_{4},H_{5}\}$ is a linearly independent set. Otherwise, $|M|=1$ and $H_{1}$ is a block-constant matrix that belongs to the space spanned by $\{H_{2},H_{3},H_{4},H_{5}\}$, concluding the proof. ∎
@@ 4.3 $k$-Densest Subgraph
Consider a graph $G=(V,E)$ exactly as in the Max-Cut problem and an integer $k$ such that $1<k<n$. The $k$-Densest Subgraph is a constrained problem that seeks the subgraph $G^{\prime}$ with exact $k$ vertices that contains the most number of edges [69]. Similarly to Max-Cut, we encode each solution in a computational basis state of an $n$-qubit system, where the corresponding bitstring $z=z_{1}\dots z_{n}$ indicates the selected vertices: $z_{i}=1$ if vertex $i$ is included in the subgraph $G^{\prime}$, and $z_{i}=0$ otherwise. On the other hand, unlike Max-Cut, $k$-Densest Subgraph is a constrained problem. In this case, feasible solutions are restricted to bitstrings with Hamming weight exactly $k$. Therefore, $|\mathcal{S}|=2^{n}$ and $|\mathcal{S^{\prime}}|=\binom{n}{k}$. Efficient indexing and un-indexing operators for the set of $k$-combinations are provided by Marsh and Wang [11].
@@ 5 Conclusion and Outlook
In this work, we have derived a general upper bound on the DLA dimension of QWOA. In particular, we showed that this bound scales polynomially in the number of distinct cost values of the cost function. An implication of our result is that, for optimization problems in the class $\mathsf{NPO}\text{-}\mathsf{PB}$, the DLA dimension is polynomially bounded in the input size. This structural insight allows us to derive meaningful consequences for both the expressivity and trainability of QWOA. By connecting our bound to recent analytical results on QWOA’s performance limitations, we identified complexity-theoretic conditions under which the QWOA ansatz must be overparameterized to either achieve exact optimality or reach a fixed approximation ratio. In particular, for problems outside the $\mathsf{BPPO}$ and $\mathsf{BP\text{-}APX}$ classes, overparameterization becomes necessary. Another significant outcome of our analysis is the demonstration that QWOA avoids barren plateaus for all problems in $\mathsf{NPO}\text{-}\mathsf{PB}$. This follows from combining our polynomially bounded DLA dimension with recent theoretical frameworks linking expressivity to variance scaling in variational quantum algorithms. Unlike prior work, which required asymptotic assumptions on the spectral distribution of the problem Hamiltonian, our result holds for arbitrary instances within this problem class, thus generalizing earlier arguments and providing broader assurance of QWOA’s trainability in practical settings.
@@
Several open questions remain. One important direction for future work is to derive non-trivial lower bounds on the DLA dimension of QWOA, which would help determine the tightness of our results and offer a more complete characterization of the algorithm’s expressivity. Establishing such tight bounds could enable sharper variance estimates for the loss function and yield a deeper understanding of QWOA’s optimization landscape. It is also worth exploring how expressivity, as quantified by DLA dimension, relates to generalization behavior in noisy or data-driven contexts, particularly in hybrid quantum-classical settings. Furthermore, our Lie-algebraic framework could be applied more broadly to other variational quantum algorithms, such as QAOA and adaptive ansätze, potentially leading to unified expressivity measures across variational paradigms. We hope that the simplicity and generality of our approach provide a foundation for these future directions and contribute to a deeper theoretical understanding of quantum optimization algorithms.
@@
Another promising direction is how our result can inspire the development of new ansätze. In this sense, QWOA is structurally agnostic by design, which grants generality but limits its performance to Grover-style quadratic speed-up. Motivated by the limitations this structural agnosticism imposes on mixing design, a growing trend in the literature is to investigate how incorporating problem structure into quantum algorithms can help overcome such constraints. Notable works in this direction include: Headley [70] used a statistical approach to analyze QAOA with both the original transverse field mixing and the line mixing, showing how the algorithm can capture correlations between solutions in both settings; Matwiejew and Wang [16] introduced a variant of QWOA called Quantum Walk Optimization Algorithms on Combinatorial Subsets (QWOA-CS), where the continuous-time quantum walk incorporates problem structure via an algebraic framework based on graph automorphisms; and more broadly, Matwiejew, Pye, and Wang [71] consider structural exploitation beyond combinatorial optimization problems, extending the discussion to continuous optimization settings. The results and insights presented here — particularly those concerning trainability and expressivity — suggest a potential impact on this growing line of research, which warrants further investigation. In particular, exploring how our framework can inform the design of structure-aware mixing may help uncover mechanisms that enable performance surpassing that of Grover’s algorithm. In this context of tailored ansätze designs for specific problem families, we can explore alternative approaches that incorporate problem structure, such as those incorporating nonlocal operators or time-dependent dynamics.
@@ Acknowledgments:
The authors would like to thank Andris Ambainis for valuable input on computational complexity classes. G.A.B. thanks the financial support of Coordenação de Aperfeiçoamento de Pessoal de Nível Superior (CAPES/Brazil) Grant No. 88887.950125/2024-00. D.L. gratefully acknowledges funding from QuantERA Project QOPT. This work is partially supported by the National Research Foundation, Singapore, and A*STAR under its CQT Bridging Grant and its Quantum Engineering Programme under grant NRF2021-QEP2-02-P05. S.A. thanks the financial support from FY23 SUG-PATRICK REBENTROST (WBS Code: A00098700000). R.A.M.S. thanks the financial support from the Latvian Quantum Initiative under the European Union Recovery and Resilience Facility project number 2.3.1.1.i.0/1/22/I/CFLA/001. F.L.M. thanks the financial support of Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq/Brazil) Grant No. 407296/2021-2.

### v2

@@ 3.1 Bounding the DLA Dimension for QWOA
We provide an upper bound on the DLA dimension for QWOA that scales quadratically in $m$. Recall that $m$ is the number of distinct cost values over the feasible solutions, or equivalently, the number of distinct eigenvalues of $H_{C}$ restricted to the feasible subspace. The key idea behind the proof is the following. Since QWOA is invariant under permutations of basis states, solutions with the same cost can be grouped in the problem Hamiltonian $H_{C}$. Due to the diagonal structure of the problem Hamiltonian $H_{C}$ and the uniform structure of the mixing Hamiltonian $H_{M}$, the nested commutators of the QWOA generators will share the same partition structure of $m^{2}$ blocks, where each block has constant entries. This means that the degree of freedom in the DLA space generated by the QWOA ansatz is bounded by $\mathcal{O}(m^{2})$.
@@
For any instance of a combinatorial optimization problem, the DLA dimension of QWOA satisfies $\operatorname{dim}(\mathfrak{g}_{\operatorname{QWOA}})\leq m^{2}+1$. In particular, $\operatorname{dim}(\mathfrak{g}_{\operatorname{QWOA}})\leq m^{2}$ when, for each $1\leq j\leq m$, the cost value $x_{j}$ either has multiplicity 11 ($k_{j}=1$) or is 0 ($x_{j}=0$).
@@
We first consider the unconstrained case. The problem Hamiltonian $H_{C}$ is diagonal in the computational basis and can be rewritten in a block diagonal form as
$H_{C}=\begin{bmatrix}x_{1}I_{k_{1}}&\bar{0}&\cdots&\bar{0}\\ \bar{0}&x_{2}I_{k_{2}}&\cdots&\bar{0}\\ \vdots&\vdots&\ddots&\vdots\\ \bar{0}&\bar{0}&\cdots&x_{m}I_{k_{m}}\\ \end{bmatrix}$, (6)
where $I_{x}$ is the identity matrix with the dimension $x$ and $\bar{x}$ denotes a matrix with all elements equal to $x$. We call such a partitioning of $H_{C}$ into blocks a block partitioning pattern. The generator $H_{M}$ can be written using the block partitioning pattern as
@@
Let $\Lambda$ be the set of skew-Hermitian matrices with the same block partitioning pattern as $H_{C}$ and $H_{M}$ such that all elements in any block are the same. In this case, note that $iH_{M}\in\Lambda$, but (in general) $iH_{C}\notin\Lambda$. From the definition of matrix multiplication and the fact that the skew-Hermitian operators are closed under the commutator, it follows that $iH_{C},A]\in\Lambda$ and $A,B]\in\Lambda$ for $A,B\in\Lambda$. Consequently, all nested commutators are elements of $\Lambda$. Observe that $\Lambda$ is isomorphic to the unitary Lie algebra $\mathfrak{u}(m)$, therefore their size is $|\Lambda|=m^{2}$. Considering the worst case, where $iH_{C}\notin\Lambda$, we can conclude that $\operatorname{dim}(\mathfrak{g}_{\operatorname{QWOA}})\leq m^{2}+1$. On the other hand, when $iH_{C}\in\Lambda$, we can tighten the bound to $\operatorname{dim}(\mathfrak{g}_{\operatorname{QWOA}})\leq m^{2}$. This case occurs if and only if, for each $1\leq j\leq m$, $k_{j}=1$ or $x_{j}=0$, finishing the unconstrained case.
@@ 3.2 Overparametrization in QWOA for NPO-PB problem
We now present two remarks regarding item 1 of Theorem 2. While these remarks are stated in the context of item 1, analogous reasoning applies to item 2 by replacing $\mathsf{BPPO}$ and $\mathsf{PO}$ with $\mathsf{BP\text{-}APX}$ and $\mathsf{APX}$, respectively. Firstly, the inverse of item 1 of Theorem 2 does not hold, i.e, $\mathcal{P}\in\textsf{BPPO}$ does not necessarily imply that the optimal QWOA depth is polynomially bounded in the input size. Indeed, the fact that a problem belongs to BPPO does not guarantee that it can be efficiently solved by random sampling, which is the baseline performance that QWOA quadratically improves upon. Secondly, restricting the theorem to $\mathsf{PO}$ rather than $\mathsf{BPPO}$ would yield a stronger result. However, unless $\mathsf{PO}=\mathsf{BPPO}$, the arguments used in the proof of item 1 do not allow us to conclude that $\mathcal{P}\notin\mathsf{PO}$ implies $p^{*}_{\mathcal{P}}\notin\mathcal{O}(\operatorname{poly}(s))$. Whether there exists a problem in $\textsf{BPP}-\textsf{PO}$ that QWOA can still solve with polynomial depth remains an open question.
@@ 4.1 Unstructured Search
$H_{1}=iH_{C}=i\begin{pmatrix}I&\overline{0}\\ \overline{0}&\overline{0}\\ \end{pmatrix},\qquad H_{2}=iH_{M}=i\begin{pmatrix}\overline{1}&\overline{1}\\ \overline{1}&\overline{1}\\ \end{pmatrix}$. (9)
@@
$H_{3}=-\begin{pmatrix}\overline{0}&\overline{1}\\ \overline{-1}&\overline{0}\\ \end{pmatrix},\qquad H_{4}=-i\begin{pmatrix}\overline{0}&\overline{1}\\ \overline{1}&\overline{0}\end{pmatrix},\qquad H_{5}=-i\begin{pmatrix}\overline{2(|M|-|S)}&\overline{2|M|-|S|}\\ \overline{2|M|-|S|}&\overline{2|M|}\\ \end{pmatrix}$. (10)
If $|M|>1$, one can show that $\{H_{1},H_{2},H_{3},H_{4},H_{5}\}$ is a linearly independent set, reaching the tight bound of Theorem 1 with $\operatorname{dim}(\mathfrak{g}_{\operatorname{QWOA,Search}})=5$. Otherwise, $|M|=1$ and Theorem 1 implies that $\operatorname{dim}(\mathfrak{g}_{\operatorname{QWOA,Search}})\leq 4$. Here, the tight bound is achieved because $H_{1}$ is a block-constant matrix that belongs to the space spanned by $\{H_{2},H_{3},H_{4},H_{5}\}$. ∎
@@ 4.3 $k$-Densest Subgraph
Consider a graph $G=(V,E)$ exactly as in the Max-Cut problem and an integer $k$ such that $1<k<n$. The $k$-Densest Subgraph is a constrained problem that seeks the subgraph $G^{\prime}$ with exactly $k$ vertices that contains the most edges [64]. Similarly to Max-Cut, we encode each solution in a computational basis state of an $n$-qubit system, where the corresponding bitstring $z=z_{1}\dots z_{n}$ indicates the selected vertices: $z_{i}=1$ if vertex $i$ is included in the subgraph $G^{\prime}$, and $z_{i}=0$ otherwise. On the other hand, unlike Max-Cut, $k$-Densest Subgraph is a constrained problem. In this case, feasible solutions are restricted to bitstrings with Hamming weight exactly $k$. Therefore, $|\mathcal{S}|=2^{n}$ and $|\mathcal{S^{\prime}}|=\binom{n}{k}$. Efficient indexing and un-indexing operators for the set of $k$-combinations are provided by Marsh and Wang [11].
@@ 5 Conclusion and Outlook
In this work, we have derived a general upper bound on the DLA dimension of QWOA. In particular, we showed that this bound scales polynomially in the number of distinct cost values of the cost function. An implication of our result is that, for optimization problems in the class $\mathsf{NPO}\text{-}\mathsf{PB}$, the DLA dimension is polynomially bounded in the input size. This structural insight allows us to derive meaningful consequences for the expressivity of QWOA. By connecting our bound to recent analytical results on QWOA’s performance limitations, we identified complexity-theoretic conditions under which the QWOA ansatz must be overparameterized to either achieve exact optimality or reach a fixed approximation ratio. In particular, for problems outside the $\mathsf{BPPO}$ and $\mathsf{BP\text{-}APX}$ classes, overparameterization — often of exponential magnitude — becomes necessary. This stands in sharp contrast with QAOA, where numerical evidence indicates that optimal solutions can be obtained with high probability as the ansatz becomes only slightly overparameterized [22]. Our result, therefore, suggests that the DLA dimension cannot always be taken as a reliable indicator of convergence in VQAs.
@@
Several open questions remain. One important path for future work is to derive non-trivial lower bounds on the DLA dimension of QWOA, which would help determine the tightness of our results and offer a more complete characterization of the algorithm’s expressivity. It is also worth exploring how expressivity, as quantified by DLA dimension, relates to generalization behavior in noisy or data-driven contexts, particularly in hybrid quantum-classical settings. Another promising avenue concerns the classical simulability of QWOA. Specifically, the results of Goh et al., combined with our Corollary 1, suggest that the expectation value of QWOA could be computed in polynomial time for $\mathsf{NPO}\text{-}\mathsf{PB}$ problems, provided certain prerequisites are met, such as the knowledge of a basis for the DLA. This potential could lead to classical simulation methods competitive with the approach of Golden et al. [65]. Furthermore, our Lie-algebraic framework could be applied more broadly to other variational quantum algorithms, such as QAOA and adaptive ansätze, potentially leading to unified expressivity measures across variational paradigms. We hope that the simplicity and generality of our approach provide a foundation for these future directions and contribute to a deeper theoretical understanding of quantum optimization algorithms.
@@
Another promising direction is how our result can inspire the development of new ansätze. In this sense, QWOA is structurally agnostic by design, which grants generality but limits its performance to Grover-style quadratic speed-up. Motivated by the limitations this structural agnosticism imposes on mixing design, a growing trend in the literature is to investigate how incorporating problem structure into quantum algorithms can help overcome such constraints. Notable works in this direction include: Headley [66] used a statistical approach to analyze QAOA with both the original transverse field mixing and the line mixing, showing how the algorithm can capture correlations between solutions in both settings; Matwiejew and Wang [16] introduced a variant of QWOA called Quantum Walk Optimization Algorithms on Combinatorial Subsets (QWOA-CS), where the continuous-time quantum walk incorporates problem structure via an algebraic framework based on graph automorphisms; and more broadly, Matwiejew, Pye, and Wang [67] consider structural exploitation beyond combinatorial optimization problems, extending the discussion to continuous optimization settings. The results and insights presented here suggest a potential impact on this growing line of research, which warrants further investigation. In particular, exploring how our framework can inform the design of structure-aware mixing may help uncover mechanisms that enable performance surpassing that of Grover’s algorithm. In this context of tailored ansätze designs for specific problem families, we can explore alternative approaches that incorporate problem structure, such as those incorporating nonlocal operators or time-dependent dynamics.
@@ Acknowledgments:
The authors would like to thank Andris Ambainis for valuable input on computational complexity classes. G.A.B. thanks the financial support of Coordenação de Aperfeiçoamento de Pessoal de Nível Superior (CAPES/Brazil) Grants No. 88887.950125/2024-00 and 88881.128522/2025-01. D.L. gratefully acknowledges funding from QuantERA Project QOPT. L.P. acknowledges support by the National Research Foundation, Singapore, and A*STAR under its CQT Bridging Grant and its Quantum Engineering Programme under grant NRF2021-QEP2-02-P05. R.A.M.S. thanks the financial support from the Latvian Quantum Initiative under the European Union Recovery and Resilience Facility project number 2.3.1.1.i.0/1/22/I/CFLA/001. F.L.M. thanks the financial support of Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq/Brazil) Grants No. 407296/2021-2 and 306049/2025-2. S.A. thanks the financial support from FY23 SUG-PATRICK REBENTROST (WBS Code: A00098700000).